OpenStudy (anonymous):

Need help ! I'm struggling so much. I will give a medal and fan anyone who helps.

OpenStudy (anonymous):

Choose the equation of the line passing through the point (-2, 4) and perpendicular to y = -negative one halfx + 4.

OpenStudy (tkhunny):

You need the slope of the line you seek. Do you know how to get that slope from the perpendicular line?

OpenStudy (anonymous):

no

OpenStudy (tkhunny):

Well, you should know that or you cannot solve this problem. With one exception, multiply two slopes and achieve -1 and they are perpendicular. m1 = 2, m2 = 3, then m1*m2 = 2*3 = 6 and they are NOT perpendicular. m1 = 1/3, m2 = 3, then m1*m2 = (1/3)*3 = 1 and the are NOT perpendicular. m1 = -1/2, m2 = 2, then m1*m2 = (-1/2)*2 = -1 and the ARE perpendicular. Can you know construct the requested equation?

OpenStudy (anonymous):

@texaschic101 @johnweldon1993